What is the force on an object of 50 kg that is 80 m away from another object that is 150 kg?

 

An object is 300m from another object of mass 40kg and experiences a force of 50N. What is the mass of the first object?

 


1
Expert's answer
2021-02-25T11:29:16-0500

Answer

Gravitational force between two masses is given

"F=\frac{Gmm'}{r^2}"

According to question

First case

"F_1=\frac{6.67×10^{−11}\times50\times150}{(80)^2}\\=7.80\times10^{−11} N"


Now mass of another object


"m=\frac{Fr^2}{Gm'}\\=\frac{50\times300^2}{6.67\times10^{-11}\times40}\\=16867\times10^{11}Kg"